Sep 30, 2011 (UT)

Data is in /newdata

modsDisp to display the newest spectra.

Observers: Fuyan Bian, christopher Willmer, betsy Green, Nathan Smith

Support Astronomer: Olga Kuhn

TO: Geno

0108 Thick clouds in the west, nice sunset...

0215 The sky is clear, we are opening.

0222 Restart the GCS,

0229 Send preset to the telescope, move to bright star to check pointing.

0235 Pointing check is done, move to a Persion star to collimate the telescope

0240 The collimation is done (<400nm). It takes less than 2 min to conver

0241 Move to the standard star bd28. Collimate the telescope. Acquisition and take spectra.

0302 Move to the first science target J2121 A. (Conditions: thin clouds in the sky and the image quality is 0.8-0.9)

0310 The script is failed (timeout error), Olga checked the script but nothing seems wrong, she notice that the mask is out (good) but the filter does not change to sdss_r (bad) in the control GUI.

0317 Resend the preset, but the AGW does not work. Restart GCS.

0323 Olga take an on sky image to see the position of the mask, and mask is not in.

0325 Try the acquisition script again. Meet the same guider problem: fail to receive full guide camera readout. Geno make a call and they will restart the machine to see whether to fix that problem.

0335 Olga rerun the script in the instrument part, it works very well. We are waiting for the guider to be fixed.

0338 Geno reset the SCAM G server computer for the guide.

0345 Finally figure out which machines to restart (SCam G and MODS 1).

45mins lost due to above problems

0349 resend the preset to J2121 A (Imaging quality is 0.79")

0402 We are taking the science exposure on J2121 A (Imaging quality is 0.9-1.0 and thin clouds, actual imaging quality is 0.8. There is about 0.1" difference between the FWHM measured by the guide and the actual imaging quality.).

0424 The first spectrum is out. The spectrum looks good. The imaging quality is getting worse. the clouds are being built up.blue is 0004 and red us 0011

0445 The second spectrum is readout. looks good. Seeing getting worse, 1.4". blue is 0005 and red is 0012

0507 The third spectrum is readout. signal is weaker due to the bad seeing. the seeing is really bad during the fourth spectrum which is 1.8". The wind speed is 10-15 m/s 006 - 0013

0528 the last two exposure is very bad finally we lost the guide star at 5:43

0543 slew to the next target for c04_131_517 to identify LBGs. (in this condition we can not get high S/N spectrum of the LBGs). can not see a 13th magnitude star from guider, change to a 11th magnitude star.

0609 There are Rain drops. I am closing.

-- FuyanBian - 29 Sep 2011.
Topic revision: r2 - 30 Sep 2011, FuyanBian
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback